OpenStudy (usukidoll):

Let T be the set of all matrices of the form AB-BA where A and B are nxn matrices. Show that span T is not Mnn. (Hint: use properties of trace)

OpenStudy (usukidoll):

Given: T is the set of all matrices with the form of AB-BA. A and B are nxn matrices. Proof: We need to prove that span T isn't Mnn. However, we need to obtain information from section 1.3 exercise 43. From Section 1.3 Exercise 43 Tr(AB)=Tr(BA) Let AB = c Tr(AB) = Tr(C) C=[Cij]  TR(AB)=TR(BA) TR(AB)-TR(BA)=0 Span T is a subset of the set S of all nxn matrices with trace = 0. However, S is a proper subset of Mnn.